OpenStudy (anonymous):

(-2)+5(x-3)-4x(6+1)a.assosiative property.b.commutative propetry. c.distributive property.d.identity property.whats the right one?

OpenStudy (anonymous):

thank you for coming

OpenStudy (anonymous):

it is supposed to pick one of the choices?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i have 5min left so what do you think?

OpenStudy (anonymous):

im devided between assossiative and commutative

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

@Mertsj please. He only has 4min left

OpenStudy (mertsj):

The problem isn't all there. There is no equal sign.

OpenStudy (anonymous):

their was no equal sigh.

OpenStudy (mertsj):

What do the directions say?

OpenStudy (anonymous):

this is weird... It has so many terms.. Each of them has a diferent property

OpenStudy (mertsj):

Put distributive. That is the property that is indicated twice.

OpenStudy (anonymous):

wich of the fallowing properies is nessasery to simplify the expression.

OpenStudy (mertsj):

There is no change in order or pairing so that rules out commutative and associative

OpenStudy (mertsj):

Distributive for sure.

OpenStudy (mertsj):

And the next time you post a question, post all of it.

OpenStudy (anonymous):

what about this question (2*6)3=2(6*3)

OpenStudy (anonymous):

-23x-17 cant be commutative or assossiative. Go for mertsj. Put distributive

OpenStudy (anonymous):

same derections.

OpenStudy (anonymous):

that one is commutative

OpenStudy (anonymous):

the directions say witch property is used?

OpenStudy (anonymous):

their is no commutative.

OpenStudy (mertsj):

associative for (2*6)3=2(6*3)

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

jeez. Im out. Need to rest /m/

OpenStudy (anonymous):

i passed the test thanks guys. :)